OpenStudy (bloomlocke367):

Solve this in interval form, from 0 to 2pi. Round to the nearest hundredth. \(-3sin2\theta=1.5\)

OpenStudy (bloomlocke367):

@mathmate @pooja195

OpenStudy (bloomlocke367):

@dan815 Will you come help?

OpenStudy (bloomlocke367):

can you help?

OpenStudy (bloomlocke367):

@pooja195

OpenStudy (bloomlocke367):

@ganeshie8 Can you please help?

ganeshie8 (ganeshie8):

as a start, divide -3 both sides

ganeshie8 (ganeshie8):

\[-3\sin(2\theta) = 1.5 \implies \sin(2\theta) = \frac{1.5}{-3} =-\frac{1}{2} \]

ganeshie8 (ganeshie8):

\[\sin(2\theta) = -\frac{1}{2} \implies 2\theta = \sin^{-1}\left(-\frac{1}{2}\right)\] look up unit circle

OpenStudy (bloomlocke367):

I have the unit circle

OpenStudy (bloomlocke367):

what am I looking for?

ganeshie8 (ganeshie8):

see for what angles the sin value is \(-\frac{1}{2}\)

OpenStudy (bloomlocke367):

sin is the y-value, right?

ganeshie8 (ganeshie8):



OpenStudy (bloomlocke367):

so, 7pi/6 and 11pi/6

OpenStudy (bloomlocke367):

and I just round that to the nearest hundredth and I'll have my answer?

ganeshie8 (ganeshie8):

Yes, \[2\theta = \frac{7\pi}{6}+2k\pi,~ ~\frac{11\pi}{6}+2k\pi\] that means \[\theta = \frac{7\pi}{12}+k\pi,~ ~\frac{11\pi}{12}+k\pi\]

ganeshie8 (ganeshie8):

but you want solutions in the interval 0 to 2pi, right ?

OpenStudy (bloomlocke367):

yes

ganeshie8 (ganeshie8):

plugin k = 0, 1, 2 etc till you hit 2pi

OpenStudy (bloomlocke367):

why?

ganeshie8 (ganeshie8):

\[\theta = \frac{7\pi}{12}+k\pi,~ ~\frac{11\pi}{12}+k\pi\] plugin \(k=0\) and get \[\theta = \frac{7\pi}{12},~ ~\frac{11\pi}{12}\] plugin \(k=1\) and get \[\theta = \frac{7\pi}{12}+\pi,~ ~\frac{11\pi}{12}+\pi\]

ganeshie8 (ganeshie8):

because any integer value of \(k\) gives an angle that is a solution to the given equation

ganeshie8 (ganeshie8):

there are infinitely many solutions as there are infinitely many integers your question is about finding the solutions in the interval 0 to 2pi

OpenStudy (bloomlocke367):

okay.

OpenStudy (bloomlocke367):

and you plug in k=2, and get \(\large\frac{7\pi}{12}+2\pi,~\frac{11\pi}{12}+2\pi\)

ganeshie8 (ganeshie8):

yes clearly thats more than \(2\pi\), so we may stop now.

ganeshie8 (ganeshie8):

k=0 and k=1 are enough

ganeshie8 (ganeshie8):

find the corresponding \(\theta\) values as you can see there are exactly four solutions in the interval 0 to 2pi

OpenStudy (bloomlocke367):

okay, and I just round it?

ganeshie8 (ganeshie8):

the question asks you to round it so yeah

OpenStudy (bloomlocke367):

so, 1.83, 2.88, 4.97, and 6.02

ganeshie8 (ganeshie8):

Looks good, you may double check with wolfram http://www.wolframalpha.com/input/?i=solve%20-3sin(2%5Ctheta)%3D1.5%2C%200%3C%5Ctheta%3C2pi&t=crmtb01

OpenStudy (bloomlocke367):

Okay, thanks!

ganeshie8 (ganeshie8):

yw